Circular No. 6593

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 970228 T. L. Cline, P. S. Butterworth, and D. E. Stilwell, Goddard Space

Flight Center, NASA; and E. P. Mazets, S. V. Golenetskii, R. L. Aptekar, D. D. Frederiks, V. N. Il’inskii, and M. M. Terekhov, Ioffe Politechnical Institute, St. Petersburg, for the Wind/Konus Gamma-Ray-Burst Team; and K. Hurley, Space Sciences Laboratory, Berkeley, for the Ulysses γ-Ray-Burst Team, report: “An independent source locus of the Feb. 28 transient event, based on observations from the interplanetary network (IPN) incorporating GGS-Wind and Ulysses, and entirely independent of observations from the BeppoSAX spacecraft, consists of a source annulus with half-width < 30′′. This γ-ray burst, first localized by BeppoSAX, has been speculated as being possibly associated with one or more candidate sources in other wavelength regions, including a long-lived x-ray component, a weak VLA radio component, and most recently an optical transient. The center of the IPN annulus is at α = 11^{h}03^{m}47^{s}, δ = +36^{o}27′.8 (equinox 2000.0) and its radius is 83^{o}.4313. The discrepancy with the SAX LECS source is < 1′, within the combined error estimates, and the discrepancy with the optical transient source is < 15′′, entirely consistent with the IPN source accuracy.”

A. Yoshida, N. Kawai, C. Otani, and F. Tokanai, Institute of Physical and Chemical Research (RIKEN); H. Inoue, T. Murakami, F. Nagase, R. Shibata, and Y. Ueda, Institute of Space and Astronautical Science; S. Holt, Goddard Space Flight Center, NASA, on behalf of the ASCA team; L. Piro and E. Costa, Istituto di Astrofisica Spaziale, CNR, Frascati; J. Heise, Space Research Organization of the Netherlands, Utrecht, on behalf of the SAX team; and J. Hughes, Rutgers University, report: “ASCA observed the SAX burst position of GRB 970228 (IAUC 6572) on Mar. 7.028–7.486 UT with a net exposure time of 20 000 s. The data displayed a weak excess in both GIS and SIS images of the revised BeppoSAX error region (IAUC 6576). Image fitting was done to both the GIS and SIS data sets, assuming a point source. The source’s location was found to be consistent with that of SAX J0501.7+1146, considering statistical errors and the attitude uncertainty of ASCA. The 2–10-keV flux and its 1σ statistical error are estimated to be (9.0 ± 2.6) × 10^{−14} erg cm^{−2} s^{−1} for the GIS and (7.2 ± 2.1) × 10^{−14} erg cm^{−2} s^{−1} for the SIS, assuming a power-law spectrum with a photon index of 1.4.”

1997 March 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green